OpenStudy (anonymous):

|3x| + 36 > 12 SOLVE INEQUALITY

OpenStudy (anonymous):

Our aim is to have x (or whatever the variable is) on its own on the left of the inequality sign: 3x +36 > 12

OpenStudy (anonymous):

3x+36-36 > 12-36 sol- 3x>-24

OpenStudy (anonymous):

-8 @Rohangrr

OpenStudy (anonymous):

nope the answer is 3x>-24

geerky42 (geerky42):

Subtract both sides by 36. \[|3x| > -24\] \(|3x|\) is always \(\ge 0\) so \(x \in \mathbb{R}\)

OpenStudy (anonymous):

THESE ARE THE ANSWERS A.-8 > x > 8 B.all real numbers are solutions C.x < -8 or x > 8 D.-8 < x < 8

OpenStudy (anonymous):

@Rohangrr

geerky42 (geerky42):

Read my reply above here. It's B.

OpenStudy (anonymous):

OK

geerky42 (geerky42):

You seem like you don't understand, do I need to clarify something?

OpenStudy (anonymous):

NO I SAW YOUR REPLY

geerky42 (geerky42):

Ok, so is it clear?

OpenStudy (anonymous):

@reaven I had given the same answer.

OpenStudy (anonymous):

YES

geerky42 (geerky42):

@Rohangrr Didn't you notice the absolute value?

OpenStudy (anonymous):

YES

geerky42 (geerky42):

Well, Rohangrr forgot to include the absolute value so his answer is wrong. just saying.

OpenStudy (anonymous):

SO MY ANSWER ISNT b @geerky42

geerky42 (geerky42):

It is B. Rohangrr's answer is different.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Remember, an absolute value is always positive. So, despite our what our algebra tells us, the absolute value has to be at least zero. After we have that in mind, we get\[\left| 3x \right|>0\]Which indicates x can be any number except zero. The solution set would look like\[x \in \mathbb{R}-\left\{ 0 \right\}=(-\infty , 0) \cup (0, \infty) = \left\{ x \in \mathbb{R}|x \ne 0 \right\}\]

OpenStudy (anonymous):

Correction to above: an absolute value is always nonnegative.

geerky42 (geerky42):

@AnimalAin but when x = 0, 36 > 12, which is true so 0 is also solution.

OpenStudy (anonymous):

Actually, I suppose my deletion of zero isn't really needed in this case.

OpenStudy (anonymous):

Solution:\[x \in \mathbb{R}\]

geerky42 (geerky42):

Yup.